Project

General

Profile

Nightsynopsistemplate

*Observing Team* _DES Observers:_ Jim Annis, Liz Buckley-Geer, Saint Nick _Telops:_ Leonardo Paredes, Javier Rojas _Remote Support:_ Klaus Honscheid, Santiago Serrano, Juan Estrada, Tom Diehl, Freddy Mercury *Observing Plan* # Guider testing # Standards at twilight, twilight # gr on deep SN fields C3, X3 # iz on deep SN field X3 # Image miniSurvey and maxVis fields *Conditions* Seeing was 0.9-1.25". Humidity was high and threatened to close us more than once. *Accomplishments* # i on deep SN field X3 # some r,i,z,y imaging in miniSurvey # r,i,z,y imaging in SPT-E/maxVis area # sunrise twilight standards: maxVis_grizy, E2a, E4a *Narrative* # The first few images were contaminated by light from an alignment laser aimed from the Cass cage up to the f/8 counterweight that was left on. # Guider testing made some progress. # The dewpoint inside the crates neared the temperature of the cooling lines, however, and we shut the dome, called on remote support and started diagnosing. Dry nitrogen isn't being sent to creates 1,2,3, it seems. Javier was advised to raise the coolant temperature for tonight, and we went back onto the sky. DECam and CTIO will meet tomorrow to decide on the next steps. # We worked X3, but it became apparent during the z exposures that the telescope was oscillating: the seeing was a few tenths worse but the ellipticity was factors of 3 worse (QR). We decided to stop observing these long exposures. This was at the equator and 2 hours past meridian. # We turned on obsTac and let it go. It mostly did maxVis, but decided to grab images in miniSurvey, as well. Surprisingly, it did some work in r, even though the moon was > 50%, though > 75 degrees away from field. Not incorrect, just surprising. # Throughout these observations at times the telescope would oscillate. The evidence for this was the guider error signals oscillating in y, the seeing getting worse, and the ellipticity > 0.15. There are graphs of the guider error signals in the logbook. At times QR would report >2" images; these were doubled. IH and kenttools reported ~1" fwhm for these images. I don't know what is right here, but the QR values are more useful for data quality and for obsTac. We believe that all images with ellipticty > 0.15 are doubled and are un-reducible. # Clearly this didn't happen last night. # On pointing: the mystery offset last night was due to the pointing model being set up to place the target in the center of N4. We left our offset in place (the TCS GUI in fact does retain information about the zeropoints) and we tracked the offsets from our target pointings. The pointing were, to first order, pretty good. At closer inspection there are systematic drifts, probably as a function of Hour Angle and Dec. The logbook has a plot. We should consider devoting more time to telescope pointing model/engineering. *SN Deadman Status* SN Deadman State E1: 2 days E2: 2 days X1: 2 days X2: 2 days X3: gr=3 day, i=4 days, z= 2 S1: 3 day S2: 5 day C1: 1 day C2: 1 day C3: 3 day *Issues* # See above # Nightsum crashes now that IH is allowed to report NULL as a seeing value. It seems to Dan Kelson and me that this may be the primary shifting. The size of the shift in the focal plane is ~5 pixels = 75 microns. Also attached is a screen shot of the motion of the primary as measured after the counterweight system was fixed. If so, the donut dx and dy should point in the direction of the gravity vector. So should the direction of shift in the guider. The fact that it is bigger in x than y, then, is an artifact us pointing in the south west; at the end of the night we'll try the north east and seeas Dan has standards there.